Two electric charges of 3 nC and -5 nC are not located at the base of an equilateral triangle (1m each side). calculate the total electric potential at the center (point P) of the triangle 1 m 1 m 1 m OA -435 OB 20.5 V OC -534 V OD 31.2 v OE-11.4V OF 24 3 V OG -15V OH-5025
Answer #1

Soln im 5 Go Inc Im - ne < 0.5 x €0860-0.5 = 1/3mi Potential at P VP = K. 3x09 KESHO 1113 1183 Yp = kxro? [3-5] 1183 Vp = -2x